All Systems Operational
Past Incidents
Feb 21, 2017

No incidents reported today.

Feb 20, 2017

No incidents reported.

Feb 19, 2017

No incidents reported.

Feb 18, 2017

No incidents reported.

Feb 17, 2017
Resolved - All emails have been processed into tickets.
Feb 17, 18:54 GMT
Monitoring - The problem is solved, no mail was lost. The backlog of emails is being processed and expects to complete in roughly 1 hour.
Feb 17, 17:30 GMT
Identified - Ticket generation from email is currently failing, we are investigating. Please use the portal for now (https://portal.pulsant.com)
Feb 17, 15:47 GMT
Resolved - As the circuits have been confirmed as stable for over 12 hours, we are now marking this incident as closed.
Feb 17, 16:22 GMT
Monitoring - All repair work is finished and circuits recovered. No further service interruptions expected.
Feb 17, 04:19 GMT
Update - Further update on progress

1.Cable preparation has started and should be completed by 01:30.
2.Splicing of the priority tubes will start from 01:30 and circuits will begin to return live.
3.Priority circuits should be complete by approximately 02:30.
4.The remainder of the joint will be spliced tube by tube straight through the joint and should be completed in the early hours of this morning.
Feb 17, 01:37 GMT
Update - Works are still progressing on this fault below is an estimated time line of works to be carried out.

**Please note this is an estimate and is subject to change as work progresses if any obstacles arise.

1. Cable will be laid by 00:01
2. Cables to be prepped.
3. Start splicing through the priority circuits around 02:00 to 03:00.
4. ETR will be given at 02:30 when we have a better understanding of what work is left to complete.
Feb 16, 22:26 GMT
Update - We have no further update at this time. We are continuing to chase our supplier on a regular basis
Feb 16, 20:06 GMT
Update - The fibre cabling has been maliciously cut and requires specialist replacement and repair work. Additional cabling crews were on standby and have been called to site with specialist equipment. Planning is underway to remove and replace the damaged cabling before the fibres are spliced into active service

We will continue to to communicate with the supplier and provide updates.
Feb 16, 19:14 GMT
Update - While carrying out repairs, the Fibre Team discovered two (2) further possible fault locations.

Both locations have been investigated and this has been reduced to one (1), fault location.

The Fibre Team have created a plan to replace a section of cable and all teams are on site carrying out work.

No ETA has been provided at this time.

We will continue to to communicate with the supplier and provide updates.
Feb 16, 18:09 GMT
Update - Engineers have opened the chamber and are currently repairing the circuit, and carrying out testing. No ETA for resolution has been provided yet.
Feb 16, 17:16 GMT
Update - Local authorities are on site and have set up traffic management to allow them to inspect the junction box and further locate the fault.

Our fibre provider expects an update within an hour from local investigations at which point they expect to have a fault resolution plan and estimated time to fix.
Feb 16, 16:23 GMT
Update - We have no further update at this time. We are continuing to chase our supplier on a regular basis.

The next update will be provided at 16:00.
Feb 16, 14:07 GMT
Update - Our supplier has confirmed that the fibre break is located beneath a major trunk road.

At this time the supplier is liaising with local authorities and traffic management to gain access to allow repair work to commence.

A further update will be provided by 14:00 unless we have additional updates before then.
Feb 16, 11:20 GMT
Update - Our supplier has advised that traffic management will be required to facilitate the fibre repair. As such, we are unable to provide an ETA for restoration of service.

A further update will be provided by 11:30am.
Feb 16, 10:21 GMT
Identified - A resilient core network link at our Milton Keynes facility is down at present due to a fibre break.

As a result, services at our Milton Keynes facility should be considered "at risk" until this is resolved.

We are working closely with our supplier to monitor the progress of repair work.

A further update will be provided at 10:30am.

If you have any questions or concerns, please raise a ticket via support@pulsant.com or by calling our Support team on 0845 11 99 999.

Please accept our apologies for any inconvenience this issue may be causing.

Kind regards,
Pulsant Support
Feb 16, 08:27 GMT
Completed - The scheduled maintenance has been completed.
Feb 17, 06:00 GMT
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Feb 16, 22:00 GMT
Scheduled - Dear Customer,

Please be aware of the following planned maintenance window for UPS Maintenance at our Maidenhead Facility, Pulsant 2 and 3 - 'B' Feed only.

WHAT
In keeping with the manufacturer's recommendations and the Pulsant planned and preventative maintenance regime, we will be carrying out scheduled UPS maintenance and updating the firmware on all UPS units at our Maidenhead Facility, Pulsant 2/3 only.

SERVICE IMPACT
NONE - There will be no planned service impact, this is an advisory note, as there will be reduced UPS resilience whilst the maintenance is being carried out.
Power will remain unaffected to customer racks, but with reduced resiliency during the maintenance period. The maintenance has been split over 2 separate weeks to ensure that the update of firmware on site has no issues. The firmware update has been confirmed by our UPS manufacturer and is required to be carried out to comply with the maintenance regime on these units.
However to bring the unit(s) back on line there will be an element of switching that needs to be done hence an "At Risk" notification is being sent out.
The maintenance will be completed by a specialist 3rd party and the UPS units will be fully tested prior to being returned to normal service. Complete power protection will be restored upon completion of the UPS maintenance.

WHY
Pulsant follow a stringent Planned Preventative Maintenance regime. This work forms a major part of this regime and the firmware update ensures full availability and continued functionality of the units in line with the manufacturers recommendations.
Pulsant's specialist UPS contractors will be on site to complete the scheduled maintenance, supervised by the datacentre management team throughout.

WHEN
to mitigate risk as far as possible and ensure all maintenance works are completed and the firmware upgrade is done, work has been scheduled to be carried out during the following maintenance window:

B Feed Only
Thursday 16th of February 2017, commencing at 22:00 GMT and ending by 06:00 GMT on Friday February 17th 2017.
Reduced resilience on the B feed. The A feed will remain unaffected and continuing to provide full resilient protection.
The "At Risk" window will be kept to an absolute minimum and the UPS systems will be returned to full resilience at the earliest opportunity.
A completed notification will be sent out once the works have been completed and normal service has resumed.
If you have any questions, please raise a ticket by emailing support@pulsant.com or by calling our support team directly.

Regards,
Pulsant Support Team
0845 1199 999
Feb 15, 14:46 GMT
Feb 15, 2017

No incidents reported.

Feb 14, 2017
Resolved - This issue has now been resolved.
Feb 14, 16:59 GMT
Identified - Issued detected with sites hosted by 123-reg
Feb 14, 10:08 GMT
Feb 13, 2017

No incidents reported.

Feb 12, 2017

No incidents reported.

Feb 11, 2017

No incidents reported.

Feb 10, 2017

No incidents reported.

Feb 9, 2017

No incidents reported.

Feb 8, 2017

No incidents reported.

Feb 7, 2017

No incidents reported.